# Regularized Langevin Dynamics for Combinatorial Optimization

### v1

@@ 3.2 Regularized Langevin Simulated Annealing
Since the gradient of the energy function could be computed in an closed form for various CO problems, here we first assume $\nabla H(\mathbf{x})$ is available. Note that the score function of the EBM could be written as $s_{\tau}(\mathbf{x})=\log p_{\tau}(\mathbf{x})=-\frac{1}{\tau}\nabla H(\mathbf{x})$. To avoid the clutter, we denote $\Delta=(2\mathbf{x}-1)\odot\nabla H(\mathbf{x})$, whose $i$-th coordinate approximates the drop of the energy function if we flip the value of $\mathbf{x}_{i}$.
Explicitly solving Equation 10 is challenging due to the presence of the sigmoid function. However, when $\tau\to 0$, the sigmoid function is approximately an indicator function:
$\lim_{\tau\to 0}\texttt{Sigmoid}(\frac{1}{2\tau}\Delta_{i}-\frac{1}{2\alpha})=\mathbbm{1}(\frac{1}{2\tau}\Delta_{i}-\frac{1}{2\alpha}>0).$ (11)
This property allows us to efficiently regularize the SA algorithm with the $d$-th largest element in $\Delta$, denoted as $\Delta_{(d)}$. We then obtain the flipping probability by letting $\frac{1}{\alpha}=\frac{\Delta_{(d)}}{\tau}$:
$q(\mathbf{x}^{\prime}_{i}=1-\mathbf{x}_{i}|\mathbf{x})=\texttt{Sigmoid}(\frac{1}{2\tau}(\Delta_{i}-\Delta_{(d)})).$ (12)
In our experiment, we find simply ignoring $-\frac{1}{\alpha}$ and normalizing all sigmoid function outputs to sum to $d$ could also work, followed by clipping all values into the range $[0,1]$. While in this work we just stick to using the $d$-th largest element. We call the resultant SA algorithm as Regularized Langevin Simulated Annealing (RLSA), whose details are summarized in Algorithm 1.
@@
4:     $\tau\leftarrow\tau_{0}(1-\frac{t-1}{T})$
@@
7:        $p\leftarrow\texttt{Sigmoid}(\frac{1}{2\tau}(\Delta_{i}-\Delta_{(d)}))$
@@
In our implementation, the elementwise sampling is run in parallel and we maintain $K$ independent SA processes simultaneously. The whole algorithm could be implemented in a few lines and accelerated with GPU-based deep learning frameworks, such as PyTorch (Paszke et al., 2017) and Jax (Bradbury et al., 2018). In this work, we implement our algorithm mainly based on PyTorch Geometric (Fey & Lenssen, 2019), and an example code is attached in Appendix C.
Given the overall RLSA framework, we now address the question of how to compute the gradient of the energy function. Numerous classical CO problems are defined over graphs and could be formulated in quadratic form, known as QUBO (Lucas, 2014). Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph, with node set $\mathcal{V}=\{1,\cdots,N\}$, edge set $\mathcal{E}\in\mathcal{V}\times\mathcal{V}$, and adjacency matrix $\mathbf{A}\in\{0,1\}^{N\times N}$. In this work, we focus on the following three problems, which have been commonly used in benchmark evaluations for CO solvers.
@@
The maximum independent set (MIS) problem aims to select the largest subset of nodes from the graph $\mathcal{G}$, without any adjacent pair. Denote a selected node as $\mathbf{x}_{i}=1$ and an unselected one as $\mathbf{x}_{i}=1$, the energy function of MIS could be expressed as
$\begin{split}H(\mathbf{x})=&-\sum_{i=1}^{N}\mathbf{c}_{i}\mathbf{x}_{i}+\beta\sum_{(i,j)\in\mathcal{E}}\mathbf{x}_{i}\mathbf{x}_{j}\\ &=-\mathbf{c}^{\top}\mathbf{x}_{i}+\beta\frac{\mathbf{x}^{\top}\mathbf{A}\mathbf{x}}{2},\end{split}$ (13)
where $\mathbf{c}\in\mathbb{R}_{+}^{N}$ is the node weight vector. It is not hard to compute the gradient of the energy function as
$\nabla H(\mathbf{x})=-\mathbf{c}+\beta\mathbf{A}\mathbf{x}.$ (14)
@@
$H(\mathbf{x})=-\sum_{i=1}^{N}\mathbf{c}_{i}\mathbf{x}_{i}+\beta\sum_{(i,j)\notin\mathcal{E}}\mathbf{x}_{i}\mathbf{x}_{j}.$ (15)
In order to represent the energy function with the adjacency matrix $\mathbf{A}$, we rewrite the penalty as $[(\sum_{i=1}^{N}\mathbf{x}_{i})^{2}-\sum_{i=1}^{N}\mathbf{x}_{i}^{2}]/2-\sum_{(i,j)\in\mathcal{E}}\mathbf{x}_{i}\mathbf{x}_{j}$, resulting in the energy function and its gradient:
$H(\mathbf{x})$ $=\mathbf{c}^{\top}\mathbf{x}+\beta\frac{(\mathbf{1}^{\top}\mathbf{x})^{2}-\mathbf{x}^{\top}\mathbf{x}-\mathbf{x}^{\top}\mathbf{A}\mathbf{x})}{2},$ (16)
$\nabla H(\mathbf{x})$ $=-\mathbf{c}+\beta((\mathbf{1}^{\top}\mathbf{x})\mathbf{1}-\mathbf{x}-\mathbf{A}\mathbf{x}).$ (17)
Max Cut.
The max cut (MCut) problem looks to partition the nodes into two sets so that the number of edges between two sets is maximized. Here we use $\mathbf{x}_{i}=1$ and $\mathbf{x}_{i}=0$ to represent the belonging to two sets, and the energy function could be expressed as
@@ 3.3 Regularized Langevin Neural Network
When the gradient is intractable or a better approximation of the locally-informed proposal in Equation 8 is wanted, we could parameterize the sampling distribution $q_{\theta}(\mathbf{x}^{\prime}|\mathbf{x})$ with a NN. Here we still utilize a mean-filed decomposition, letting $q_{\theta}(\mathbf{x}^{\prime}|\mathbf{x})=\Pi_{i=1}^{N}q_{\theta}(\mathbf{x}^{\prime}_{i}|\mathbf{x})$. The RLD update in Equation 9 could be translated into the following training loss
$\begin{split}l_{RLD}(\theta;\mathbf{x},d,\lambda)=&\mathbf{E}_{q_{\theta}(\mathbf{x}^{\prime}|\mathbf{x})}[H(\mathbf{x}^{\prime})]\\ +&\lambda(\sum_{i}^{N}q_{\theta}(\mathbf{x}^{\prime}_{i}=1-\mathbf{x}_{i}|\mathbf{x})-d)^{2}.\end{split}$ (20)
Here the first term is similar to the loss function used in the Erdoes Goes Neural (Karalias & Loukas, 2020b), which maximizes the conditional expectation of the energy function after the update. While the second term regularizes the expected Hamming distance between the two solutions, with $\lambda$ being the regularization coefficiednt. We name this NN-based solver as Regularized Langevin Neural Network (RLNN).
We train RLNN in a similar fashion to reinforcement learning through sampling and update, but without the need to account the future states except the immediate next one. This allows RLNN to circumvent the high variance in estimating the future return when trained with a long sampling process. In detail, each time we sequentially samples $T^{\prime}$ samples with the current proposal distribution $q_{\theta}(\mathbf{x}^{\prime}|\mathbf{x})$, then for each sample, we train RLNN to minimize the loss in Equation 20. The training algorithm is summarized in Algorithm 2.
Algorithm 2 Regularized Langevin Neural Network (train)
@@ 3.4 Connection to Normalized Gradient Descent
NGD is developed to address the vanishing/exploding gradient by normalizing the L2222 norm of the gradient for a scale-invariant update at each step. Our method could be treated as a discrete version of method by restricting the L2222 distance between the solutions before and after the update.
@@
The key difference between the two lies in the case when $\Delta_{(d)}<0$, RLD could not be translated into a gradient descent algorithm in Equation 4 (in terms of minimizing the energy function) since $\alpha=\frac{\tau}{\Delta_{(d)}}<0$ reverses the gradient descent direction. Instead, this should be treated as a way to escape the local optima without dramatically increasing the energy function. As an analogy n to Equation 5, we can express this situation as
$q(\mathbf{x}^{\prime}|\mathbf{x})=\frac{\exp(-\frac{\Delta_{(d)}}{2\tau}\|\mathbf{x}^{\prime}-\mathbf{x}+\frac{1}{2\Delta_{(d)}}H(\mathbf{\mathbf{x}})\|_{2}^{2})}{Z(\mathbf{x})}.$ (22)
@@
The density of $q(\mathbf{x}^{\prime}|\mathbf{x})$ increases with respect to the distance from $\mathbf{x}-\frac{1}{2\Delta_{(d)}}H(\mathbf{\mathbf{x}})$ , which is the gradient ascent direction (note $\Delta_{(d)}$ is negative here) of the energy function. This behavior arises due to the different property of the local optima in the discrete data, which may not vanish to $\mathbf{0}$ but point to an infeasible region (with $\Delta$ negative in all dimensions).
@@
Let us take MIS as an example, whose local optima is a maximal independent set, i.e., each unselected node has at least one neighbor in the set. The gradient for the selected node ($\mathbf{x}_{i}=1$) is $\nabla H(\mathbf{x})_{i}=-\mathbf{c}_{i}<0$, pointing to the direction of increasing the value, which is infeasible as $\mathbf{x}_{i}\leq 1$. While the gradient for an unselected node ($\mathbf{x}_{i}=0$) is lower bounded by $\nabla H(\mathbf{x})_{i}\geq-\mathbf{c}_{i}+\beta>0$, pointing to the direction of decreasing the value, which is also infeasible. Since the gradient descent direction is not informative, RLD would try to escape this local optima but avoid the steepest direction to increase the energy function, which is exactly the gradient ascent direction: $\mathbf{x}-\frac{1}{2\Delta_{(d)}}H(\mathbf{\mathbf{x}})$.
@@
With the same example, we could also see why the standard discrete Langevin sampler (Zhang et al., 2022) with a constant step size does work well. Since LD is a first-order approximation of the locally-informed proposal (Zanella, 2017), a small $\alpha$ is needed to make the approximation accurate. However, a small $\alpha$ would also lead to a strong penalization on the update. While at the local optima, we also have $\Delta_{i}<0$ discourage the change, as indicated in Equation 7. Therefore, additional efforts are needed to help LD escape the local optima beyond the force of random noise. Such a distinction between CO and continuous optimization highlights the significance of RLD.
@@ 4 Experiments
Table 1: Comparative results on the Mximum independent Set (MIS) problem. On each dataset, we bold the best result and color the second-best one in green. By ”best” or ”second best”, we exclude the OR solvers (Gurobi and KaMIS) as their running times are excessively large, preventing a fair comparison with the methods in other categories.
@@
Table 2: Comparative results on the Max Clique (MCl) and Max Cut (MCut) problems. On each dataset, we bold the best result and color the second-best one in green. By ”best” or ”second best”, we exclude the OR solvers (Gurobi and KaMIS) as their running times are excessively large, preventing a fair comparison with the methods in other categories.

### v2

@@ 4.2 Regularized Langevin Simulated Annealing
When $\mathbf{x}$ is binary, we could follow Zhang et al. (2022) to let $k(\mathbf{x},\mathbf{x}^{\prime})=\exp\Bigl{(}-\frac{\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}^{2}}{2\alpha}\Bigr{)}$ and explicitly write out the expectation in Equation 10 with the flipping probabilities:
$\sum_{i=1}^{N}\sigma\biggl{(}\frac{1}{2}s(\mathbf{x})_{i}(1-2\mathbf{x}_{i})-\frac{1}{2\alpha}\biggr{)}=d.$ (11)
Since the gradient of the energy function could be computed in a closed form for various CO problems, here we first assume $\nabla H(\mathbf{x})$ is available. Note that the score function of the EBM could be written as
$s_{\tau}(\mathbf{x})=\log p_{\tau}(\mathbf{x})=-\frac{1}{\tau}\nabla H(\mathbf{x}).$ (12)
To avoid clutter, we denote $\Delta=(2\mathbf{x}-1)\odot\nabla H(\mathbf{x})$, whose $i$-th coordinate approximates the drop of the energy function if we flip the value of $\mathbf{x}_{i}$.
The exact solving of Equation 11 is challenging due to the presence of the sigmoid function. However, when $\tau\to 0$, we observe that the sigmoid function is approximately an indicator function:
$\lim_{\tau\to 0}\sigma\biggl{(}\frac{1}{2\tau}\Delta_{i}-\frac{1}{2\alpha}\biggr{)}=\mathbbm{1}\biggl{(}\frac{1}{2\tau}\Delta_{i}-\frac{1}{2\alpha}>0\biggr{)}.$ (13)
This property allows us to efficiently regularize the SA algorithm with the $d$-th largest element in $\Delta$, denoted as $\Delta_{(d)}$. We then obtain the flipping probabilities by letting $\frac{1}{\alpha}=\frac{\Delta_{(d)}+\epsilon}{\tau}$, where $\epsilon\geq 0$ (e.g., $10^{-6}$) is optional:
$q(\mathbf{x}^{\prime}_{i}=1-\mathbf{x}_{i}|\mathbf{x})=\sigma\biggl{(}\frac{1}{2\tau}(\Delta_{i}-\Delta_{(d)}-\epsilon)\biggr{)}.$ (14)
We call the resultant SA algorithm as Regularized Langevin Simulated Annealing (RLSA), whose details are summarized in Algorithm 1.
@@
4:     $\tau\leftarrow\tau_{0}\bigl{(}1-\frac{t-1}{T}\bigr{)}$
@@
7:        $p\leftarrow\sigma\bigl{(}\frac{1}{2\tau}(\Delta_{i}-\Delta_{(d)}-\epsilon)\bigr{)}$
@@
It is worth noting that $k(\mathbf{x},\mathbf{x}^{\prime})=\exp\Bigl{(}-\frac{\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}^{2}}{2\alpha}\Bigr{)}$ is chosen here just for simplicity, but other strategies may perform just as well. For instance, one could omit the $-\frac{1}{\alpha}$ term and instead normalize the sigmoid outputs:
$\tilde{p}_{i}$ $=\frac{\sigma\bigl{(}\tfrac{1}{2}\,s(\mathbf{x})_{i}\,(1-2\mathbf{x}_{i})\bigr{)}}{\sum_{j=1}^{N}\sigma\bigl{(}\tfrac{1}{2}\,s(\mathbf{x})_{j}\,(1-2\mathbf{x}_{j})\bigr{)}}\,d,$ (15)
$p_{i}$ $=\min\{1,\max\{0,\tilde{p}_{i}\}\},$ (16)
so that $\sum_{i}p_{i}\approx d$. Investigating alternative kernel functions or normalization strategies—potentially adapted to specific problem structures—offers a promising avenue for future work aimed at further minimizing the approximation error.
In our implementation, the elementwise sampling is run in parallel and we maintain $K$ independent SA processes simultaneously. The whole algorithm could be implemented in a few lines and accelerated with GPU-based deep learning frameworks, such as PyTorch (Paszke et al., 2017) and Jax (Bradbury et al., 2018). An example PyTorch code is attached in Appendix C.
Given the overall RLSA framework, we now address the question of how to compute the gradient of the energy function. Numerous CO problems are defined on graphs and could be formulated in the quadratic form, known as QUBO (Lucas, 2014). Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph, with node set $\mathcal{V}=\{1,\cdots,N\}$, edge set $\mathcal{E}\in\mathcal{V}\times\mathcal{V}$, and adjacency matrix $\mathbf{A}\in\{0,1\}^{N\times N}$. In this work, we focus on the following three problems, which have been commonly used in benchmark evaluations for CO solvers.
@@
The maximum independent set (MIS) problem aims to select the largest subset of nodes of the graph $\mathcal{G}$, without any adjacent pair. Denote a selected node as $\mathbf{x}_{i}=1$ and an unselected one as $\mathbf{x}_{i}=0$, the energy function of MIS could be expressed as
$\begin{split}H(\mathbf{x})=&-\sum_{i=1}^{N}\mathbf{x}_{i}+\beta\sum_{(i,j)\in\mathcal{E}}\mathbf{x}_{i}\mathbf{x}_{j}\\ &=-\mathbf{1}^{\top}\mathbf{x}+\beta\frac{\mathbf{x}^{\top}\mathbf{A}\mathbf{x}}{2},\end{split}$ (17)
As this is a quadratic function, it is straightforward to compute the gradient of the energy function as
$\nabla H(\mathbf{x})=-\mathbf{1}+\beta\mathbf{A}\mathbf{x}.$ (18)
@@
$H(\mathbf{x})=-\sum_{i=1}^{N}\mathbf{x}_{i}+\beta\sum_{(i,j)\notin\mathcal{E}}\mathbf{x}_{i}\mathbf{x}_{j}.$ (19)
In order to represent the energy function with the adjacency matrix $\mathbf{A}$, we can write the penalty as $\Bigl{(}\bigl{(}\sum_{i=1}^{N}\mathbf{x}_{i}\bigr{)}^{2}-\sum_{i=1}^{N}\mathbf{x}_{i}^{2}\Bigr{)}/2-\sum_{(i,j)\in\mathcal{E}}\mathbf{x}_{i}\mathbf{x}_{j}$, reformulating the energy function and its gradient as
$H(\mathbf{x})$ $=-\mathbf{1}^{\top}\mathbf{x}+\beta\frac{(\mathbf{1}^{\top}\mathbf{x})^{2}-\mathbf{x}^{\top}\mathbf{x}-\mathbf{x}^{\top}\mathbf{A}\mathbf{x}}{2},$ (20)
$\nabla H(\mathbf{x})$ $=-\mathbf{1}+\beta\bigl{(}(\mathbf{1}^{\top}\mathbf{x})\mathbf{1}-\mathbf{x}-\mathbf{A}\mathbf{x}\bigr{)}.$ (21)
Maximum Cut.
The maximum cut (MCut) problem looks to partition the nodes into two sets so that the number of edges between two sets is maximized. Here we use $\mathbf{x}_{i}=1$ and $\mathbf{x}_{i}=0$ to represent the belonging to two sets, and the energy function could be expressed as
@@ 4.3 Regularized Langevin Neural Network
The efficacy of gradient-guided SA solvers, including RLSA, is critically dependent on the knowledge of $\nabla H(\mathbf{x})$ (Sun et al., 2022). However, $\nabla H(\mathbf{x})$ often lacks a closed-form expression or is too costly to compute directly, necessitating the gradient approximation. To evaluate RLD under these conditions, we employ a neural network to approximate the sampling distribution $q_{\theta}(\mathbf{x}^{\prime}|\mathbf{x})$, thereby demonstrating its effectiveness even when only an approximate gradient is available.
Here we still utilize a mean-field decomposition, letting $q_{\theta}(\mathbf{x}^{\prime}|\mathbf{x})=\Pi_{i=1}^{N}q_{\theta}(\mathbf{x}^{\prime}_{i}|\mathbf{x})$. The RLD update in Equation 10 could be translated into the following training loss
$\begin{split}l_{RLD}&(\theta;\mathbf{x},d,\lambda)=\mathbb{E}_{q_{\theta}(\mathbf{x}^{\prime}|\mathbf{x})}[H(\mathbf{x}^{\prime})]\\ +&\lambda\biggl{(}\sum_{i}^{N}q_{\theta}(\mathbf{x}^{\prime}_{i}=1-\mathbf{x}_{i}|\mathbf{x})-d\biggr{)}^{2}.\end{split}$ (24)
The first term minimizes the conditional expectation of the energy function after the one-step update. When the expectation is tractable, this term is equivalent to the unsupervised learning loss of Erdőes Goes Neural (EGN) (Karalias & Loukas, 2020); otherwise, we could optimize this loss by estimating the policy gradient, i.e., $\mathbb{E}_{q_{\theta}(\mathbf{x}^{\prime}|\mathbf{x})}[H(\mathbf{x}^{\prime})\nabla\log q_{\theta}(\mathbf{x}^{\prime}|\mathbf{x})]$, via Monte Carlo methods. In this work, we focus on verifying the effectiveness of RLD and adopts the unsupervised training loss for simplicity. The second term regularizes the expected Hamming distance between the two solutions, with $\lambda$ being the regularization coefficient. We name this NN-based solver as Regularized Langevin Neural Network (RLNN).
We train RLNN in a similar fashion to reinforcement learning through sampling and update, but without the need to account the future states except the immediate next one. This allows RLNN to circumvent the high variance in estimating the future return when trained with a long sampling process. In detail, each time we sequentially sample $T^{\prime}$ samples with the current proposal distribution $q_{\theta}(\mathbf{x}^{\prime}|\mathbf{x})$, then for each sample, we train RLNN to minimize the loss in Equation 24. The training algorithm of RLNN is summarized in Algorithm 2.
Algorithm 2 Regularized Langevin Neural Network
@@ 4.4 Connection to Normalized Gradient Descent
NGD is developed to address the vanishing/exploding gradient by normalizing the L2222 norm of the gradient for a scale-invariant update at each step. Our method, especially RLSA, could be treated as a discrete version of this approach by regularizing the Hamming distance between the solutions before and after the update. And both methods accelerate the descent under the smoothness condition.
@@
The key difference between the two lies in the case when $\Delta_{(d)}<0$, RLSA could not be translated into a gradient descent algorithm to minimize the energy function, since $\alpha=\frac{\tau}{\Delta_{(d)}}<0$ (ignore $\epsilon$) reverses the direction of gradient descent. Instead, RLSA should be treated now as a way to escape local optima without dramatically increasing the energy function. Utilizing Equation 5, we can fully express the proposal at this state as
$q(\mathbf{x}^{\prime}|\mathbf{x})=\frac{\exp\bigl{(}-\frac{\Delta_{(d)}}{2\tau}\|\mathbf{x}^{\prime}-\mathbf{x}+\frac{1}{2\Delta_{(d)}}\nabla H(\mathbf{x})\|_{2}^{2}\bigr{)}}{Z(\mathbf{x})}.$ (26)
@@
It should be noted that the density of $q(\mathbf{x}^{\prime}|\mathbf{x})$ increases with respect to the distance from $\mathbf{x}-\frac{1}{2\Delta_{(d)}}\nabla H(\mathbf{\mathbf{x}})$, which is the gradient ascent direction (note that $\Delta_{(d)}$ is negative here) of the energy function. This behavior is desirable due to the different property of the local optima in the discrete data, which may not vanish to zero but point to an infeasible region (i.e., with $\Delta$ negative in all coordiantes).
@@
Let us take MIS as an example, whose local optimum corresponds to a maximal independent set, that is, each unselected node has at least one neighbor in the set. At the local optimum, the gradient at the selected node $\mathbf{x}_{i}$ is $\nabla H(\mathbf{x})_{i}=-1$, which points to the direction of increasing $\mathbf{x}_{i}$, and is infeasible since $\mathbf{x}_{i}\leq 1$. Similarly, the gradient at an unselected node is bounded by $\nabla H(\mathbf{x})_{i}\geq-1+\beta>0$, which points to the direction of decreasing the value, and is also infeasible. Since the gradient descent direction is not informative, RLSA would try to escape this local optima but avoid the steepest direction to increase the energy function, i.e., the gradient ascent direction $\mathbf{x}-\frac{1}{2\Delta_{(d)}}\nabla H(\mathbf{\mathbf{x}})$.
@@
With the same example, we could also see why the standard discrete Langevin sampler (Zhang et al., 2022) with a constant step size fails here. Since LD is a first-order approximation of the locally-balanced proposal (Zanella, 2017), a small $\alpha$ is needed to make the approximation accurate. However, a small $\alpha$ would also lead to a strong penalization on the magnitude of the update. At local optima, $\Delta_{i}<0$ would further discourage the change, as indicated in Equation 7. Therefore, additional efforts are needed to help LD escape the local optima beyond the force of random noise. This distinction between combinatorial and continuous optimization highlights the significance of RLD.
@@ 5 Experiments
Table 1: Comparative results on the Mximum independent Set (MIS) problem. On each dataset, we bold the best result and color the second-best one in green. By “best” or “second best”, we exclude the OR solvers (Gurobi and KaMIS) as their runtime is excessively large, preventing a fair comparison with the methods in other categories.
@@
Table 2: Comparative results on the Maximum Clique (MCl) and Maximum Cut (MCut) problems. On each dataset, we bold the best result and color the second-best one in green. By “best” or “second best”, we exclude the OR solvers (Gurobi and SDP) as their runtime is excessively large, preventing a fair comparison with the methods in other categories.